OpenStudy (anonymous):

verify cos(a+b)cos(a-b) = cos^2b-sin^a

OpenStudy (anonymous):

do you want to PROVE the sum formula?

myininaya (myininaya):

http://openstudy.com/users/enlightened#/updates/534ad228e4b0dfc5ac71df10 looks like you asked that question here and freckles gave you great advice

OpenStudy (anonymous):

and difference formula*

OpenStudy (anonymous):

I think the distributive property answer is really simple

myininaya (myininaya):

i am going to ask you the same question freckles asked can you give me what cos(a+b) is equal to and what cos(a-b) is equal to?

OpenStudy (anonymous):

Hint: cos(a+b) = cos(a+b) because of the property x =x

OpenStudy (anonymous):

LOL I think he means in terms of cos(a) and cos(b) and sin(a) and sin(b)

myininaya (myininaya):

@Enlightened if you want my help can you please answer my question?

OpenStudy (anonymous):

oh i'm sorry, yeah it's in terms of cos a and cos b and sina and sinb, there are no actual values to them.

OpenStudy (anonymous):

...those are the terms, you must relate them.

myininaya (myininaya):

@enlightened will you answer my question. I'm trying to get you to use the sum and difference identities on cos(a+b) and cos(a-b) you have not done that yet.

OpenStudy (anonymous):

cos(a+b)=cos(a)cos(b)-sin(a)sin(b) cos(a-b)=cos(a)cos(b)+sin(a)sin(b)

OpenStudy (anonymous):

@myininaya there it is ^

myininaya (myininaya):

great now multiply those

OpenStudy (anonymous):

what do i multiply?

myininaya (myininaya):

cos(a+b)cos(a-b) multiply the expansions for these that you gave me

myininaya (myininaya):

this is what freckles was talking about (c-d)(c+d)=c^2-d^2

OpenStudy (anonymous):

I don't understand, where does c-d,c+d come from and do use distributive to multiply all those terms?

myininaya (myininaya):

multiply cos(a+b)cos(a-b) by using the expansions you gave me

myininaya (myininaya):

it would be helpful to memorize (c-d)(c+d)=c^2-d^2 though

myininaya (myininaya):

you gave me this cos(a+b)=cos(a)cos(b)-sin(a)sin(b) cos(a-b)=cos(a)cos(b)+sin(a)sin(b) all I'm asking you to do right now is multiply (cos(a)cos(b)-sin(a)sin(b))(cos(a)cos(b)+sin(a)sin(b))

myininaya (myininaya):

this is in the form (c-d)(c+d) which equals c^2-d^2

myininaya (myininaya):

where c= and d=?